# SocialJax: An Evaluation Suite for Multi-agent Reinforcement Learning in Sequential Social Dilemmas

### v1

@@ Abstract
Social dilemmas pose a significant challenge in the field of multi-agent reinforcement learning (MARL). Melting Pot is an extensive framework designed to evaluate social dilemma environments, providing an evaluation protocol that measures generalization to new social partners across various test scenarios. However, running reinforcement learning algorithms in the official Melting Pot environments demands substantial computational resources. In this paper, we introduce SocialJax, a suite of sequential social dilemma environments implemented in JAX. JAX is a high-performance numerical computing library for Python that enables significant improvements in the operational efficiency of SocialJax on GPUs and TPUs. Our experiments demonstrate that the training pipeline of SocialJax achieves a 50× speedup in real-time performance compared to Melting Pot’s RLlib baselines. Additionally, we validate the effectiveness of baseline algorithms within the SocialJax environments. Finally, we use Schelling diagrams to verify the social dilemma properties of these environments, ensuring they accurately capture the dynamics of social dilemmas. Our project is available at https://sites.google.com/view/socialjax.
Machine Learning, ICML
@@ 1 Introduction
Solving sequential social dilemmas remains a pivotal challenge in multi-agent reinforcement learning (MARL). Melting Pot (Leibo et al., 2021) and Melting Pot 2.0 (Agapiou et al., 2022) provide comprehensive environments for testing and evaluating a broad range of research topics, such as social dilemmas, reciprocity, resource sharing, and task partitioning. Conducting experiments on Melting Pot is highly challenging, as it requires a large number of environment time steps for training. Consequently, using Melting Pot for experiments requires either extremely high hardware specifications or a substantial amount of computing time. Melting Pot is designed under the assumption that general-purpose hardware (CPU) is inexpensive, whereas specialized accelerator hardware (GPU/TPU) is costly. Under this assumption, the computational efficiency ratio between CPU and TPU using Melting Pot is approximately 100:1. Although JAX (Bradbury et al., 2018) can provide some acceleration for algorithms in Melting Pot, it cannot be used to speed up environment execution since Melting Pot environments run on CPUs. Even if both Melting Pot environments and learning algorithms were implemented in JAX, data would still need to transmit over the network between machines, because Melting Pot was originally designed around remote inference, where the CPUs running the environment are not co-located with the TPUs that are performing learning.
In Melting Pot (Leibo et al., 2021), A3C (Mnih, 2016) is used to adapt to the assumption that CPU resources are vastly more abundant than TPU/GPU resources. In Melting Pot 2.0 (Agapiou et al., 2022), algorithms that are better suited for abundant CPU resources and limited TPU/GPU resources are trained, including ACB (Espeholt et al., 2018), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020). Since both the environments and algorithms in our SocialJax implementation can run on GPUs, we can choose PPO (Schulman et al., 2017) as the baseline algorithm. Compared to the A3C/VTrace/ACB family, PPO does not require as much novel experience per update step. PPO employs a clipped loss function to prevent excessive policy updates, enhancing training stability. The A3C/VTrace/ACB family utilizes asynchronous parallelism to improve data collection efficiency, making it well-suited for distributed architectures. However, this approach increases the variance of the policy distribution, which can negatively impact training stability.
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory
@@
Figure 1: The SocialJax suite contains 6 multi-agent reinforcement learning environments designed to evaluate social dilemmas. The screenshots in this figure were taken from a third-person visualization perspective. The agents of all environments are restricted to partial observability of their surroundings through a designated observation window.
In RL research, several approaches have been proposed that leverage JAX (Bradbury et al., 2018) to execute environments on GPUs, thereby accelerating computation. PureJaxRL (Lu et al., 2022) leverages JAX to parallelize environments and executes both the algorithms and the environments on the GPU, significantly improving the efficiency of RL. However, PureJAXRL is limited to single-agent environments and does not support multi-agent scenarios. JaxMARL (Rutherford et al., 2023) extends the efficiency gains of JAX to MARL by providing a suite of JAX-based MARL environments and algorithms. While JaxMARL provides JAX-based environments for MARL, most of these environments focus on traditional cooperative settings.
To the best of our knowledge, our work is the first evaluation framework leveraging JAX for sequential social dilemmas. We build multiple social dilemma environments based on JAX, which were previously CPU-based. Moreover, we show that our work significantly accelerates the speed of environment interactions, providing greater convenience and possibilities for research on social dilemma problems.
There has been significant progress in research on sequential social dilemmas, with many studies (Hughes et al., 2018; McKee et al., 2020; Yang et al., 2020) achieving notable breakthroughs. However, one challenge they face is the lack of well-established evaluation standards. The Melting Pot (Agapiou et al., 2022; Leibo et al., 2021) framework presents some basic algorithms, such as A3C (Mnih, 2016), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020), but these algorithms are not open-sourced in the architecture of Melting Pot 2.0. Different testing environments for social dilemmas are being widely used, making it challenging to verify algorithm performance on a unified standard. SocialJax provides a different framework for evaluating algorithms targeting social dilemmas, addressing the issue of inconsistent evaluation standards in this field.
Our contributions are as follows: First, we implement six sequential social dilemma environments based on JAX and evaluate the effectiveness of IPPO (De Witt et al., 2020; Rutherford et al., 2023) under both individual and common reward settings. The environments (Figure 1), originated from Melting Pot, include Coins, Commons Harvest: Open, Commons Harvest: Closed, Clean Up, Territory, and Coop Mining, all of which feature commonly studied mixed incentives structures. Second, we show that training IPPO in SocialJax is at least 50 times faster than Melting Pot 2.0 on a machine with 14 CPUs and 1 Nvidia A100 GPU. Third, using Schelling diagrams, we further analyze the behavior of agents trained with individual and collective rewards to understand the underlying incentive structures.
@@ 2 Related Work
Sequential Social Dilemmas Libraries (Vinitsky et al., 2019) introduces a testbed for evaluating social dilemmas. However, its scope is restricted to only two environments, Clean Up and Harvest. OpenSpiel (Lanctot et al., 2019) provides a diverse set of environments, including those where cooperative and competitive dynamics coexist, making them ideal for studying social dilemmas. Melting Pot (Leibo et al., 2021) was the first to establish a standardized benchmark for MARL, specifically designed to evaluate agent performance in sequential social dilemmas. By providing a diverse set of carefully designed multi-agent scenarios, it enables systematic testing of agent behaviors in cooperative, competitive, and mixed-motive settings. Melting Pot 2.0 (Agapiou et al., 2022) introduces a more comprehensive and refined suite of testing environments that capture a wide range of interdependent relationships and incentive structures. These improvements allow researchers to better assess generalization, emergent behaviors, and strategic adaptation in MARL, making Melting Pot 2.0 a more powerful tool for studying sequential social dilemmas.
JAX-based RL environments and Algorithms JAX (Bradbury et al., 2018) is an open-source library developed by Google, designed for high-performance numerical computing and machine learning tasks. Based on the GPU and TPU acceleration capabilities introduced by JAX, along with its support for parallelism and vectorization, JAX can effectively accelerate both environments and algorithms on hardware accelerators (GPUs, TPUs). Functorch (Horace He, 2021) provides JAX-like composable function transforms for PyTorch, making it easier to accelerate RL algorithms on hardware accelerators. (Lu et al., 2022) introduces a parallelized framework leveraging JAX, enabling both environment and model training to run on GPUs.
@@
Jumanji (Bonnet et al., 2023) has developed a diverse range of single-agent environments in JAX, ranging from simple games to NP-hard combinatorial problems. Gymnax (Lange, 2022) provides JAX versions of different environments, including classic control, Bsuite (Osband et al., 2019), MinAtar (Young & Tian, 2019), and a collection of classic RL tasks. Pgx (Koyamada et al., 2024) implements a variety of board game simulation environments using JAX, such as Chess, Shogi, and Go. Brax (Freeman et al., 2021) is a physics engine written in JAX for RL, which re-implements MuJoCo. XLand-MiniGrid (Nikulin et al., 2023) is a grid-world environment and benchmark fully implemented in JAX. Similarly, our work is also demonstrated using grid-world environments. Mava (Mahjoub et al., 2024) and JaxMARL (Rutherford et al., 2023) both provide JAX-based MARL environments and baseline algorithms. Among the JAX environments and algorithms, only the Storm and Coins environments in JaxMARL are related to social dilemmas, while none of them specifically address sequential social dilemmas. We construct multiple environments focused on social dilemma challenges.
@@ 3 SocialJax
We introduce a JAX version of sequential social dilemma environments that delivers substantial performance gains through parallel processing of both algorithmic computations and environmental simulations on the GPU. Our work presents an efficient testing framework for sequential social dilemmas, complete with evaluation environments and benchmark baselines. Our SocialJax encompasses a diverse suite of environments (Coins, Commons Harvest: Open, Commons Harvest: Closed, Clean Up, Territory, and Coop Mining), which collectively capture a wide range of sequential social dilemma scenarios, including public good dilemmas and common pool resource problems. We evaluate the environments using IPPO and provided an intuitive demonstration of their social dilemma properties through $Schelling$ $diagrams$ (Perolat et al., 2017; Schelling, 1973).
@@ 3.1 Sequential Social Dilemmas
In $N$-player partially observable Markov games (Littman, 1994), agents possess partial observation of the environment. At each state, players choose actions from their respective action sets $\mathcal{A}^{1},\ldots,\mathcal{A}^{N}$. State transitions $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times\ldots\times\mathcal{A}^{N}\to\Delta(\mathcal{S})$ are determined by the current state and the joint actions taken by all agents, where $\Delta(\mathcal{S})$ represents the set of discrete probability distributions over $\mathcal{S}$. We define the observation space of player $i$ as $O_{i}=\{o^{i}|s\in\mathcal{S},o^{i}=O(s,i)\}$. Each agent has its corresponding individual reward defined as $r^{i}:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\to\mathbb{R}$. In an $N$-player $sequential$ $social$ $dilemma$, cooperators and defectors each have their own disjoint policies$(\pi_{c}^{1},\dots,\pi_{c}^{l},\dots,\pi_{d}^{1},\dots,\pi_{d}^{m})\in\Pi_{c}^{l}\times\Pi_{d}^{m}$ with $l+m=N$. $l$ is the number of cooperators. $R_{c}(l)$ and $R_{d}(l)$ are denoted as the average payoff for cooperating policies and defecting policies. When we have a sequential social dilemma, if it satisfies the following conditions (Leibo et al., 2017; Hughes et al., 2018):
1. Mutual cooperation is greater than mutual defection:
$R_{c}(N)>R_{d}(0).$
2. Mutual cooperation is greater than exploited by defectors:
$R_{c}(N)>R_{c}(0).$
3. Either the fear property or the greed property, or both:
  • Fear: mutual defection is greater than being exploited:
$R_{d}(i)>R_{c}(i)$ for sufficiently small $i$.
  • Greed: exploiting a cooperator is greater than mutual cooperation:
$R_{d}(i)>R_{c}(i)$ for sufficiently large $i$.
We use $Schelling$ $diagrams$ (Perolat et al., 2017; Schelling, 1973) to depict the payoffs of different numbers of cooperators and defectors. The lines $R_{c}(l+1)$ and $R_{d}(l)$ depict the average reward for an agent choosing to either cooperate or defect, as a function of the number of co-players that are cooperating (see Figure 3).
@@ 3.2 Environments of SocialJax
Our environments are mainly derived from Melting Pot 2.0 (Agapiou et al., 2022). The Melting Pot suite of environments is developed using an entity-component system implemented in Lua and configured via Python. This system is built on top of the DMLab2D engine, which is written in C++ and configured using Lua (Beattie et al., 2020). Since the core environment computations are always executed in C++ on the CPUs, and the algorithms run on the TPU/GPU, the overall execution speed is constrained by the efficiency of the CPUs and the data transfer bandwidth between the CPUs and GPUs. To enhance the runtime efficiency of the environments, we need to scale up the number of CPUs, which presents significant challenges to the hardware capabilities. In contrast, the environments and algorithms provided by our SocialJax implementation run entirely on GPUs, leveraging the JAX framework to significantly improve execution efficiency.
Figure 1 shows the layout of our environments. The data format observed by agents in SocialJax environments differ from those in Melting Pot 2.0 (Agapiou et al., 2022). SocialJax environments are similar to MiniGrid (Chevalier-Boisvert et al., 2018), where agents observe a grid matrix, with different objects represented by distinct numerical values. In contrast, Melting Pot 2.0 provides observations to agents in the form of image pixels. Melting Pot 2.0 was originally designed to use $8\times 8$ pixels per cell, resulting in observations of $88\times 8$ 88 pixels in most environments. In this setup, the agents have a partially observable window of $11\times 1$ 11 cells. In the Melting Pot Contest 2023 (Trivedi et al., 2024), all academic participants down-sampled the cells to $1\times 1$, also producing observations of $11\times 1$ 11. The agents in our environments have the same $11\times 1$ 11 field of view, but they are implemented at the grid level.
@@
Coins was first introduced by (Lerer & Peysakhovich, 2017) to investigate how agents behave in situations where they can cooperate with or harm others through their actions, making it a classic example of a social dilemma. Two players collect coins in a shared room, where each coin is assigned to one player’s color. Each coin, upon appearing, has an equal 50% probability of being assigned to the first player’s color or the second player’s color. An agent receives a reward of 1 for collecting any coin, regardless of its color. If one player collects a coin assigned to the other player’s color, the other player receives a reward of -2.
Commons Harvest The core concept of Commons Harvest was first defined by (Janssen et al., 2010). (Perolat et al., 2017) introduced it in multi-agent scenarios. Melting Pot 2.0 (Agapiou et al., 2022) has made it a more comprehensive environment. There are several patches of apples in the room. The players can receive a reward of 1, when they collect one apple. The apples will regrow with probability that is determined by the number of neighborhood apples in radius 2, when the apples are collected by the players. The probability of apple regrowth decreases as the nearby apples diminish. Apples will not regrow, if there are no apples in the neighborhood. Commons Harvest includes two scenarios: Commons Harvest: Closed and Commons Harvest: Open.
@@
In Commons Harvest: Closed , the players’ roles are divided into two groups: one focuses on collecting apples, while the other is tasked with preventing over-harvesting by other players (Perolat et al., 2017). There are two rooms in this setup, each containing multiple patches of apples and a single entry. These rooms can be defended by specific players to prevent others from accessing the apples by zapping them with with a beam, causing them to respawn.
@@
In Commons Harvest: Open, all individuals can harvest apples, but must restrict their actions to not collect the last apple and ensure that the apples can regrow (Perolat et al., 2017).
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory
Figure 2: Training curves of Common Reward IPPO and Individual Reward IPPO. Common Reward encourages collective interests, leading to higher overall returns, while Individual Reward primarily drives selfish behavior, often resulting in lower returns.
@@
Clean Up is a multiplayer game where players earn rewards by collecting apples, each worth +1 reward. Apples grow in an orchard and their regrowth depends on the cleanliness of a nearby river. Pollution accumulates in the river at a constant rate and once pollution surpasses a certain threshold, the apple growth rate drops to zero. Players have the option to perform a cleaning action that removes small amounts of pollution from the river. For continuous apple growth, the group must keep river pollution levels consistently low over time.
Territory is a competitive multi-agent game where players paint walls to capture territory and earn rewards. The players claim walls by painting them with their unique color, either by touching or flinging paint. The paint dries after 25 steps, brightens, and rewards are given every subsequent timestep. The more walls that are painted and dried, the greater the reward players can receive. Walls can be zapped twice to permanently destroy them, making them non-claimable and traversable. Players can also zap each other and a player is permanently removed after being hit twice.
Nine players compete to claim resources by painting them in their unique colors. Players spawn in an open area near each other, away from the resource walls. Resource distribution is uneven, with some areas contain more walls than others.
Coop Mining In this environment, two types of ore spawn randomly in empty spaces. Players can extract the ore to get reward. Iron ore (gray) can be mined individually and provides a reward of +1 upon extraction. In contrast, gold ore (yellow) requires coordinated mining by two to four players within a 3-step window, granting a reward of +8 to each participant. When a player begins to mine gold ore, the state of the ore changes to “partially mined” to indicate readiness for another player to assist. Visually, this is represented by a brighter shade of yellow. If no other player cooperates or too many players attempt to mine simultaneously, the ore reverts to its original state, and no reward is granted.
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory

### v2

@@ Abstract
Sequential social dilemmas pose a significant challenge in the field of multi-agent reinforcement learning (MARL), requiring environments that accurately reflect the tension between individual and collective interests. Previous benchmarks and environments, such as Melting Pot, provide an evaluation protocol that measures generalization to new social partners in various test scenarios. However, running reinforcement learning algorithms in traditional environments requires substantial computational resources. In this paper, we introduce SocialJax, a suite of sequential social dilemma environments and algorithms implemented in JAX. JAX is a high-performance numerical computing library for Python that enables significant improvements in operational efficiency. Our experiments demonstrate that the SocialJax training pipeline achieves at least 50× speed-up in real-time performance compared to Melting Pot’s RLlib baselines. Additionally, we validate the effectiveness of baseline algorithms within SocialJax environments. Finally, we use Schelling diagrams to verify the social dilemma properties of these environments, ensuring that they accurately capture the dynamics of social dilemmas. Our code is available at https://github.com/cooperativex/SocialJax.
@@ 1 Introduction
Solving sequential social dilemmas remains a pivotal challenge in multi-agent reinforcement learning (MARL). Efficient and rich environments and benchmarks are crucial for enabling rigorous evaluation and meaningful comparison of algorithms. Traditional MARL environments Lowe et al. (2017); Samvelyan et al. (2019); Bard et al. (2020); Carroll et al. (2019) have overwhelmingly focused on fully cooperative or competitive tasks. Only a handful of benchmarks, such as Prisoner’s Dilemma of OpenSpiel Lanctot et al. (2019), Stag Hunt and Chicken in PettingZoo Terry et al. (2021), address social dilemmas. However, these environments rely on CPU-based parallelism, which limits their simulation throughput and scalability, and they offer only a narrow variety of scenarios without standardized evaluation protocols for sequential social dilemmas.
Although there are environments Leibo et al. (2021); Agapiou et al. (2022) specifically designed for sequential social dilemmas, conducting experiments in these environments is highly challenging due to the substantial number of environment time steps required for training. Training in such environments demands extremely high hardware specifications (over 1000 CPUs for Melting Pot) or a significant amount of computing time. Moreover, while the Melting Pot framework Agapiou et al. (2022); Leibo et al. (2021) includes several baseline algorithms, such as A3C Mnih (2016), V-MPO Song et al. (2019), and OPRE Vezhnevets et al. (2020), these implementations are not publicly available, which poses challenges for reproducibility and further development.
@@
Figure 1: The SocialJax suite contains nine multi-agent reinforcement learning environments designed to evaluate social dilemmas. The screenshots in this figure were taken from a third-person visualization perspective. The agents of all environments are restricted to partial observability of their surroundings through a designated observation window.
Recent libraries such as JAX Bradbury et al. (2018) facilitate efficient execution of NumPy-style Python code on hardware accelerators(GPUs and TPUs) by automatically translating high-level functions into optimized, parallelizable operations. In reinforcement learning(RL) research, several approaches like PureJaxRL Lu et al. (2022) have been proposed to leverage JAX Bradbury et al. (2018) to run environments on GPUs. Parallelizing environments on GPUs could significantly accelerate computation. However, PureJAXRL is limited to single-agent environments and does not support multi-agent scenarios. JaxMARL (Rutherford et al., 2023) extends JAX’s efficiency gains to MARL by providing a suite of JAX-based environments and algorithms. While JaxMARL incorporates a limited set of social dilemma environments (e.g., Coins and STORM), it remains primarily geared toward traditional cooperative tasks and lacks dedicated benchmarks specifically designed for sequential social dilemmas.
In this paper, we develop a suite of environments for sequential social dilemmas based on JAX to improve computational efficiency. We also provide JAX-based implementations of multiple MARL algorithms and construct a complete training pipeline, significantly accelerating the training process. Solving sequential social dilemma problems often requires a large number of time steps and diverse evaluation environments. SocialJax addresses these challenges by offering both computationally efficient and diverse environments. Furthermore, we validate the social dilemma properties of each environment using Schelling diagrams. In addition, we design metrics beyond return for each environment, enabling a more precise quantification of the cooperative and competitive behaviors of agents. To the best of our knowledge, our work is the first evaluation framework leveraging JAX for sequential social dilemmas. Our contributions are as follows:
Implementations of Social Dilemma Environments: We implement a diverse suite of sequential social dilemma environments in JAX (see Figure 1), each exhibiting mixed-incentive dynamics, and achieve substantial improvements in environment simulation speed.
Implementation of Algorithms in JAX: We implement a series of JAX-based MARL algorithms, including Independent PPO (IPPO) (De Witt et al., 2020; Rutherford et al., 2023) under both individual and common reward settings, Multi-Agent PPO (MAPPO) (Yu et al., 2022) and Social Value Orientation (SVO) (McKee et al., 2020). We also provide training results with and without parameter sharing. We show that training algorithms in SocialJax is at least 50 times faster than Melting Pot 2.0 on a machine with 14 CPUs and 1 Nvidia A100 GPU.
Performance Benchmark and Analysis: We benchmark the performance of various environments and algorithms in terms of simulation throughput and training speed. Additionally, we evaluate the performance of the algorithms we include across various environments. At last, we design specific evaluation metrics for each environment to measure agent behaviors as either prosocial or antisocial. We further analyze the behavior of selfish and cooperative agents to understand the underlying incentive structures using Schelling diagrams Perolat et al. (2017); Schelling (1973). Finally, we analyze the performance of the implemented algorithms in our environments, including cases where agent behavior varies with different SVO McKee et al. (2020) reward angles.
@@ 2 Related Work
In this section, we review related work on sequential social dilemmas and JAX-based reinforcement learning environments, algorithms, and benchmarks. Sequential Social Dilemma Vinitsky et al. (2019) introduces a testbed for evaluating social dilemmas. However, its scope is restricted to only two environments on Clean Up and Harvest. OpenSpiel (Lanctot et al., 2019) offers a collection of environments, some of which feature both cooperative and competitive dynamics, allowing for the exploration of social dilemmas. Melting Pot Leibo et al. (2021) was the first to establish a standardized benchmark for MARL specifically designed to evaluate agent performance in sequential social dilemmas. By providing a diverse set of carefully designed multi-agent scenarios, it enables systematic testing of agent behaviors in cooperative, competitive, and mixed-motive settings. Melting Pot 2.0 Agapiou et al. (2022) introduces a more comprehensive and refined suite of testing environments that capture a wide range of interdependent relationships and incentive structures. The existing social dilemma environments are CPU-based, and their execution efficiency is constrained by CPU performance. To overcome this limitation, we developed a GPU-accelerated environment using JAX.
@@
JAX Bradbury et al. (2018) is an open-source library developed by Google, designed for high-performance numerical computing and machine learning tasks. Based on the GPU and TPU acceleration capabilities introduced by JAX, along with its support for parallelism and vectorization, JAX can effectively accelerate both environments and algorithms on hardware accelerators (GPUs, TPUs). Functorch Horace He (2021) provides JAX-like composable function transforms for PyTorch, making it easier to accelerate RL algorithms on hardware accelerators. Lu et al. (2022) introduces a parallelized framework leveraging JAX, enabling both environment and model training to run on GPUs. Jumanji Bonnet et al. (2023) has developed a diverse range of single-agent environments in JAX, ranging from simple games to NP-hard combinatorial problems. Gymnax Lange (2022) provides JAX versions of different environments, including classic control, Bsuite Osband et al. (2019), MinAtar Young & Tian (2019), and a collection of classic RL tasks. Pgx Koyamada et al. (2024) implements a variety of board game simulation environments using JAX, such as Chess, Shogi, and Go. Brax Freeman et al. (2021) is a physics engine written in JAX for RL, which re-implements MuJoCo. XLand-MiniGrid Nikulin et al. (2023) is a grid-world environment and benchmark fully implemented in JAX. Similarly, our work is also demonstrated using grid-world environments. Mava Mahjoub et al. (2024) and JaxMARL Rutherford et al. (2023) both provide JAX-based MARL environments and baseline algorithms. Among the JAX environments and algorithms, only the Storm and Coins environments in JaxMARL are related to social dilemmas, while none of them specifically address sequential social dilemmas. We construct multiple environments focused on social dilemma challenges.
@@ 3 SocialJax
In this section, we define the concept of social dilemmas, introduce the environments we implement, describe the algorithms used, and present the metrics we design to measure cooperative and competitive behaviors. Our work presents an efficient testing framework for sequential social dilemmas, complete with evaluation environments and benchmark baselines. Our SocialJax encompasses a diverse suite of environments, which collectively capture a wide range of sequential social dilemma scenarios, including public good dilemmas and common pool resource problems. We evaluate the environments using IPPO individual reward (both with and without parameter sharing), IPPO common reward, MAPPO, and SVO.
@@ 3.1 Sequential Social Dilemmas
In $N$-player partially observable Markov games Littman (1994), agents possess partial observation of the environment. At each state, players choose actions from their respective action sets $\mathcal{A}^{1},\ldots,\mathcal{A}^{N}$. State transitions $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times\ldots\times\mathcal{A}^{N}\to\Delta(\mathcal{S})$ are determined by the current state and the joint actions taken by all agents, where $\Delta(\mathcal{S})$ represents the set of discrete probability distributions over $\mathcal{S}$. We define the observation space of player $i$ as $O_{i}=\{o^{i}|s\in\mathcal{S},o^{i}=O(s,i)\}$. Each agent has its corresponding individual reward defined as $r^{i}:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\to\mathbb{R}$. In an $N$-player $sequential$ $social$ $dilemma$, cooperators and defectors each have their own disjoint policies$(\pi_{c}^{1},\dots,\pi_{c}^{l},\dots,\pi_{d}^{1},\dots,\pi_{d}^{m})\in\Pi_{c}^{l}\times\Pi_{d}^{m}$ with $l+m=N$. $l$ is the number of cooperators. $R_{c}(l)$ and $R_{d}(l)$ are denoted as the average payoff for cooperating policies and defecting policies. When we have a sequential social dilemma, if it satisfies the following conditions Leibo et al. (2017); Hughes et al. (2018):
  1. 1.
    Mutual cooperation is greater than mutual defection: $R_{c}(N)>R_{d}(0).$
  2. 2.
    Mutual cooperation is greater than being exploited by defectors:$R_{c}(N)>R_{c}(0).$
  3. 3.
    Either the fear property or the greed property (or both) holds:
    Fear: $R_{d}(i)>R_{c}(i)$ for sufficiently small $i$;  Greed: $R_{d}(i)>R_{c}(i)$ for sufficiently large $i$.
We use Schelling diagrams Perolat et al. (2017); Schelling (1973) to depict the payoffs of different numbers of cooperators and defectors. The lines $R_{c}(l+1)$ and $R_{d}(l)$ depict the average reward for an agent choosing to either cooperate or defect, as a function of the number of co-players that are cooperating (see Figure 3).
@@ 3.2 Environments of SocialJax
Figure 1 shows the layout of our environments. SocialJax environments are mainly derived from Melting Pot 2.0 Agapiou et al. (2022). However, the layout of Socialjax environments is not the same as Melting Pot, but rather more similar to MiniGrid Chevalier-Boisvert et al. (2018). Agents inside the environments observe a grid matrix, with different objects represented by distinct numerical values. In all our environments, agents have the same $11\times 1$ 11 field of view implemented at the grid level. For further details, please refer to the Appendix A. Next, we detail the environment setup, including the specific environments in our evaluation suite.
@@
Coins This environment was first introduced by Lerer & Peysakhovich (2017) to investigate how agents behave in situations where they can cooperate with or harm others by eating other agent’s type of coin, making it a classic example for social dilemma.
Commons Harvest Agents need to consume apples to obtain rewards; however, the probability of apple regrowth is determined by the number of apples in the local neighborhood. This creates tension between selfish agents and cooperative agents. In order to discuss different social behaviors of agents, in Commons Harvest, we established three different scenarios: Commons Harvest: Closed, Commons Harvest: Open, and Commons Harvest: Partnership
@@
In Commons Harvest: Closed, there are two rooms in this setup Perolat et al. (2017), each containing multiple patches of apples and a single entry. These rooms can be defended by specific players to prevent others from accessing the apples by zapping them with with a beam, causing them to respawn.
@@
In Commons Harvest: Open, all individuals can harvest apples, but must restrict their actions to not collect the last apple and ensure that the apples can regrow.
The Commons Harvest: Partnership is similar to Commons Harvest: Closed, specific players are required to defend the room. However, in this case, each room has two entry points, necessitating two players to cooperate in defending the room.
@@
Clean Up This environment is a public good game where players earn rewards by collecting apples. But the spawn rate of apples depends on the cleanliness of a nearby river. For continuous apple growth, the agents must keep river pollution levels consistently low over time.
Coop Mining In this environment, two types of ore spawn randomly in empty spaces. Iron ore (gray) can be mined individually and provides a reward of +1 upon extraction, whereas gold ore requires group coordination and yields a higher payoff per miner. Selfish agents tend to mine the iron ore more, while cooperative agents will try to cooperate and mine gold ore.
Mushrooms This environment includes multiple types of mushrooms that, when consumed by agents, deliver different rewards. Some mushrooms benefit only the individual agent, while others are beneficial to the group.
Gift Refinement In this environment, agents can give collected tokens to other agents on the map, which upgrade to yield greater rewards. Therefore, maximizing returns requires agents to trust one another and cooperate effectively.
Prisoners Dilemma: Arena Agents inside this environment collect “defect” or “cooperate” tokens and then interact with each other to compare inventories. The resulting payoffs follow the classic Prisoner’s Dilemma matrix, emphasizing the tension between individual incentives and collective welfare.